# Analytical short- and long-range kink-like structures in scalar field models with polynomial interactions

### v1

@@ Abstract
Introduction. – Kinks are topological structures that arise under the action of a single real scalar field in $(1,1)$ dimensions [1]. They are static solutions of the equation of motion which connect two distinct minima of the potential. A well-known model which supports these structures is the so-called $\phi^{4}$ model. In this case, the kink is described by a hyperbolic tangent, so its asymptotic behavior is given by an exponential function. These objects are of current interest in Physics due to their applications in distinct areas of nonlinear science. For instance, they can be used to describe specific behaviors in magnetic materials [2], to study two and three-body interaction in Bose-Einstein condensates [3, 4] and to model braneworld scenarios with a single extra dimension of infinite extent [5, 6]. Modifications in the internal structure of kinks are also of interest, since they can be used in geometrically constrained systems of magnetic domain walls [7]. In this direction, one may consider the inclusion of a parameter in the potential, as proposed in Ref. [8, 9], which modifies the slope of the kink.
@@
where $m$ is the classical mass defined in Eq. (2), and $\mu$ and $\alpha>1$ are constants which depend on the model. If $\mu=0$, i.e., the term of lowest order is linear in the field, controlled by the classical mass, so the tail has an exponential falloff, $\tilde{\phi}\propto e^{-m_{vi}|x|}$. This can be seen, for instance, in the well-known $\phi^{4}$ potential,
@@
which supports two minima, at $\phi=\pm 1$, both with classical mass $m_{\pm}^{2}=4$. In this case, $\phi(x)=\tanh(x)$, so the solution has an exponential tail, $\tilde{\phi}(x)\propto e^{-2|x|}$. There are several other potentials with this feature, including the sine-Gordon and polynomials of higher powers; see Refs. [8, 27, 14, 38].
The case in which the classical mass is null leads to a distinct tail. Since $\alpha>1$, Eq. (6) leads us to get $\tilde{\phi}(x)\propto|x|^{-2/(\alpha-1)}$, so the falloff is slower, with power-law tails, and these configurations are called long-range solutions. This occurs, for example, in the potential
@@
where $a$ is a positive real parameter. A similar model was considered in Ref. [8], but our model allows us to get a fixed maximum and classical mass modified by the parameter $a$; this leads to distinct features. The case $a=1$ recovers the potential in Eq. (7). For $10<a<$ 1, the potential supports two minima, at $\phi=\pm 1$, both of them with classical mass $m^{2}_{\pm 1}=4/a$. The point $\phi=0$ is a local maximum for $11/3\leq a<$ 1 and minimum for $30<a<1/$ 3. Interestingly, for $30<a<1/$ 3, potential presents two symmetric maxima, at $\phi=\pm\sqrt{(3a-1)/(3a-3)}$. If $a>1$, the potential engenders the aforementioned two minima at $\phi=\pm 1$ and additional others, at $\phi=\pm\sqrt{a/(a-1)}$, with infinite classical mass. The points $\phi=0$ and $\phi=\pm\sqrt{(3a-1)/(3a-3)}$ are local maxima. As $a$ gets larger and larger, the classical mass associated to the minima $\phi=\pm 1$ tends to zero and the minima with infinite classical mass run away, disappearing for $a\to\infty$. In the limit $a\to\infty$, we have
@@
In the above expression, $M(a)=-{\rm arctanh}\big{(}\sqrt{1-a}\big{)}$ for $a<1$, and $M(a)=\arctan\big{(}\sqrt{a-1}\big{)}$ otherwise. The energy is $E=4/3$ for $a=1$. In the limit $a\to\infty$, we get $E=3\pi/8$. We remark that it diverges in the limit $a\to 0$, justifying the exclusion of $a=0$. The energy is monotonically decreasing with $a$.
@@
where the $\phi_{\pm}(x)$ stands for the solution in the sector $I_{\pm}$. Notice that the solutions are now half-compact: the tails connecting the minima $\phi=\pm\sqrt{a/(a-1)}$ are compact, whilst the ones going to the minima $\phi=\pm 1$ are exponential. Interestingly, the extension of the tail which exponential falloff depends on $a$, according to $\phi_{\pm}(x)\approx\pm 1\pm(2/a)e^{-2|x|/\sqrt{a}}$ for $x\to\pm\infty$. Moreover, the parameter $a$ controls the amplitude of the solutions: as it gets larger, the amplitude gets smaller. In the limit $a\to\infty$, the above solutions become uniform, $\phi_{\pm}=\pm 1$. From Eq. (4), we get the energy density
@@
Linear Stability. – Since we are dealing with a model with a parameter that modifies the tail of the solution, let us now investigate its stability against small fluctuations. By taking $\phi(x,t)=\phi_{s}(x)+\psi(x,t)$ in Eq. (3), where $\phi_{s}(x)$ is the solution of Eq. (5) and $\psi(x,t)$ is the perturbation, we get
@@
Figure 5: The stability potential $U(x)$ in Eqs. (23), for $a=1/3,2/5,1/2,2/3$ and $1$ (left) and $a=1,5/4,3/2,2,4$ and $a\to\infty$ (right). The dotted lines represent the case $a=1$, the dash-dot line denotes the case $a=3/2$ that delimits the maximum value of $a$ for which the stability potential does not engender a volcano profile, and the dashed line stands for the limit $a\to\infty$ described by Eq. (24).
@@
The eigenvalues and number of bound states depend on the parameter $a$. For $a>1.577$, we only have a single bound state, which is the zero mode. In the range $11.577\geq a>$ 1, we have two bound states. In particular, for $a=1.577$, we have $\omega_{1}^{2}=2.536$ in addition to the zero mode, at the same height of the asymptotic values of the stability potential. For $a=1$, there are two bound state and a semi-bound state at the asymptotic value of $U(x)$, $\omega^{2}=4$ [39]. In the interval $0.4141>a>$ 0.414, there are three bound states. The case $a=0.414$ has two bound states and a semi-bound state at $\omega^{2}=9.661$. Four bound states arise if $0.1110.414>a>$ 0.111 and at least five bound states are present for $a<0.111$. In summa, as $a$ approaches zero, more and more bound states appear. In Fig. 6, we display the behavior of the bound states with $a$.
Figure 6: The eigenvalues $\omega^{2}$ in terms of the parameter $a$ in the range $(0,4]$ (left) and $(0,1]$ (right). The dashed lines represent the classical mass, $m^{2}_{\pm}=U(\pm\infty)$, associated to the stability potential in Eq. (23). The green lines represent the zero mode and the red, blue, purple and orange lines denote, respectively, the first, second, third and fourth excited states. The diamonds represent the semi-bound states.
@@
Figure 7: The finite region of the stability potential $U_{+}(x)$ (left) and $U_{-}(x)$ (right) in Eq. (27) for $a=1.75,2,2.25$ and $2.5$.
Outlook. – In this paper, we have investigated a class of polynomial potentials that allows us to increase the range of kinks with analytical results. The procedure takes advantage of the first-order framework developed by Bogomol’nyi [37] and consists of introducing a parameter in the potential that controls the classical masses.
First, we have considered the central sector $\phi\in[-1,1]$ of the potential (9) to show that, as the parameter $a$ gets larger, the solution gets a longer range, going from exponential to power-law tails. This transition also appears in the energy density. We have also investigated the linear stability to show that the parameter $a$ also modifies the shape of the stability potential, that tends to become a volcano as $a$ gets larger; this impacts the eigenvalues and bound states associated to the Schrödinger-like equation.

### v2

@@ Abstract
I Introduction
Kinks are topological structures that arise under the action of a single real scalar field in $(1,1)$ dimensions [1]. They are static solutions of the equation of motion which connect two distinct minima of the potential. A well-known model which supports these structures is the so-called $\phi^{4}$ model. In this case, the kink is described by a hyperbolic tangent, so its asymptotic behavior is given by an exponential function. These objects are of current interest in Physics due to their applications in distinct areas of nonlinear science. For instance, they can be used to describe specific behaviors in magnetic materials [2], to study two and three-body interaction in Bose-Einstein condensates [3, 4] and to model braneworld scenarios with a single extra dimension of infinite extent [5, 6]. Modifications in the internal structure of kinks are also of interest, since they can be used in geometrically constrained systems of magnetic domain walls [7]. In this direction, one may consider the inclusion of a parameter in the potential, as proposed in Ref. [8, 9], which modifies the slope of the kink.
@@ II Model
where $m$ is the classical mass defined in Eq. (2), and $\mu$ and $\alpha>1$ are constants which depend on the model. If $m\neq 0$, i.e., the term of lowest order is linear in the field, controlled by the classical mass, so the tail has an exponential falloff, $\tilde{\phi}\propto e^{-m_{vi}|x|}$. This can be seen, for instance, in the well-known $\phi^{4}$ potential,
@@
which supports two minima, at $\phi=\pm 1$, both with classical mass $m_{\pm}^{2}=4$. In this case, $\phi(x)=\tanh(x)$, so the solution has an exponential tail, $\tilde{\phi}(x)\propto e^{-2|x|}$. There are several other potentials with this feature, including the sine-Gordon and polynomials of higher powers; see Refs. [8, 27, 14, 38, 39, 40, 41, 42].
The case in which the classical mass is null ($m=0$) leads to a distinct tail. Since $\alpha>1$, Eq. (6) leads us to get $\tilde{\phi}(x)\propto|x|^{-2/(\alpha-1)}$, so the falloff is slower, with power-law tails, and these configurations are called long-range solutions. This occurs, for example, in the potential
@@
where $a$ is a positive real parameter. A similar model was considered in Ref. [8], but our model allows us to get a fixed maximum and classical mass modified by the parameter $a$; this leads to distinct features. The case $a=1$ recovers the potential in Eq. (7). For $10<a<$ 1, the potential supports two minima, at $\phi=\pm 1$, both of them with classical mass $m^{2}_{\pm 1}=4/a$. The point $\phi=0$ is a local maximum for $11/3\leq a<$ 1 and minimum for $30<a<1/$ 3. Interestingly, for $30<a<1/$ 3, potential presents two symmetric maxima, at $\phi=\pm\sqrt{(3a-1)/(3a-3)}$. If $a>1$, the potential engenders the aforementioned two minima at $\phi=\pm 1$ and additional others, at $\phi=\pm\sqrt{a/(a-1)}$, with infinite classical mass. The infiniteness of the classical mass of the latter minimum induces the compactification of its associated tail in the solution, leading to a half-compact kink. The points $\phi=0$ and $\phi=\pm\sqrt{(3a-1)/(3a-3)}$ are local maxima. For $0<a\leq 1$, there is only a single topological sector, $\phi\in[-1,1]$. For $a>1$, the presence of the aforementioned four minima allows for the existence of three topological sectors, $\phi\in[-\sqrt{a/(a-1)},-1]$, $\phi\in[-1,1]$ and $\phi\in[1,\sqrt{a/(a-1)}]$. Each topological sector supports kink and antikink solutions connecting the minima which define the sector.
As $a$ gets larger and larger, the classical mass associated to the minima $\phi=\pm 1$ tends to zero and the minima with infinite classical mass run away, disappearing for $a\to\infty$. In the limit $a\to\infty$, we have
@@
In the above expression, $M(a)=-{\rm arctanh}\big{(}\sqrt{1-a}\big{)}$ for $a<1$, and $M(a)=\arctan\big{(}\sqrt{a-1}\big{)}$ otherwise. The above expression can be expanded around $a=1$, in the form
$E=\frac{4}{3}-\frac{2}{15}(a-1)+{\cal O}\!\left[(a-1)^{2}\right]$ (17)
Therefore, the limit $a\to 1$ leads to the energy $E=4/3$, as expected, since $a=1$ recovers the so-called $\phi^{4}$ model. In the limit $a\to\infty$, we get $E=3\pi/8$. We remark that it diverges in the limit $a\to 0$, justifying the exclusion of $a=0$. The energy is monotonically decreasing with $a$.
@@
where the $\phi_{\pm}(x)$ stands for the solution in the sector $I_{\pm}$. Notice that the solutions are now half-compact: the tails associated to the minima $\phi=\pm\sqrt{a/(a-1)}$ are compact, whilst the ones going to the minima $\phi=\pm 1$ are exponential. Interestingly, the extension of the tail which exponential falloff depends on $a$, according to $\phi_{\pm}(x)\approx\pm 1\pm(2/a)e^{-2|x|/\sqrt{a}}$ for $x\to\pm\infty$. Moreover, the parameter $a$ controls the amplitude of the solutions: as it gets larger, the amplitude gets smaller. In the limit $a\to\infty$, the above solutions become uniform, $\phi_{\pm}=\pm 1$. From Eq. (4), we get the energy density
@@
III Linear Stability
Since we are dealing with a model with a parameter that modifies the tail of the solution, let us now investigate its stability against small fluctuations. By taking $\phi(x,t)=\phi_{s}(x)+\psi(x,t)$ in Eq. (3), where $\phi_{s}(x)$ is the solution of Eq. (5) and $\psi(x,t)$ is the perturbation, we get
@@ III Linear Stability
Figure 5: The stability potential $U(x)$ in Eqs. (24), for $a=1/3,2/5,1/2,2/3$ and $1$ (left) and $a=1,5/4,3/2,2,4$ and $a\to\infty$ (right). The dotted lines represent the case $a=1$, the dash-dot line denotes the case $a=3/2$ that delimits the maximum value of $a$ for which the stability potential does not engender a volcano profile, and the dashed line stands for the limit $a\to\infty$ described by Eq. (25). In the left panel, the curve with lowest asymptotic value is represented by $a=1$ (dotted line); as $U(\pm\infty)$ goes up, $a$ decreases. In the right panel, $a=1$ (dotted line) stands for the highest asymptotic value; as $U(\pm\infty)$ goes down, $a$ increases.
@@
The eigenvalues and number of bound states depend on the parameter $a$. To find them, one must use numerical procedures. For $a>1.577$, we only have a single bound state, which is the zero mode. In the range $11.577\geq a>$ 1, we have two bound states. In particular, for $a=1.577$, we have $\omega_{1}^{2}=2.536$ in addition to the zero mode, at the same height of the asymptotic values of the stability potential. For $a=1$, there are two bound state and a semi-bound state at the asymptotic value of $U(x)$, $\omega^{2}=4$ [43]. In the interval $0.4141>a>$ 0.414, there are three bound states. The case $a=0.414$ has two bound states and a semi-bound state at $\omega^{2}=9.661$. Four bound states arise if $0.1110.414>a>$ 0.111 and at least five bound states are present for $a<0.111$. In summa, as $a$ approaches zero, more and more bound states appear. In Fig. 6, we display the behavior of the bound states with $a$. It is worth commenting that the presence of the vibrational modes may lead to resonance phenomena in kink-antikink collisions. This was shown in Ref. [31], where the authors studied collisions of kinks with power-law tails and found that a rich resonance phenomenology.
Figure 6: The eigenvalues $\omega^{2}$ in terms of the parameter $a$ in the range $(0,4]$ (left) and $(0,1]$ (right). The dashed lines represent the classical mass, $m^{2}_{\pm}=U(\pm\infty)$, associated to the stability potential in Eq. (24). The green dash-dotted lines represent the zero mode. The red, blue, purple and orange lines denote, respectively, the first, second, third and fourth excited states. The thickness of the solid lines increases as one ranges from the first to the fourth excited state. The diamonds represent the semi-bound states.
@@
It is worth commenting that, even though we have shown that the solution is stable under scalar perturbations, there are other stabilities that may be of interest. For instance, one may investigate the stability against contractions and dilations. In this direction, as shown in Ref. [45], the presence of the first-order equation (11) is compatible with the stressless condition. This ensures that the solution also comply with this stability criterion. Moreover, the non-negative character of the potential (9) and the topological character of the solution (12) ensures that the Bogomol’nyi bound [37] is saturated, so the energy of the solution (12) is a global minimum of the system. Therefore, the solution cannot decay, being stable.
Figure 7: The finite region of the stability potential $U_{+}(x)$ (left) and $U_{-}(x)$ (right) in Eq. (28) for $a=1.75,2,2.25$ and $2.5$. The thickness of the lines increases with $a$.
IV Outlook
In this paper, we have investigated a class of polynomial potentials that allows us to increase the range of kinks with analytical results. The procedure takes advantage of the first-order framework developed by Bogomol’nyi [37] and consists of introducing a parameter in the potential that controls the classical masses.
First, we have considered the central sector $\phi\in[-1,1]$ of the potential (9) to show that, as the parameter $a$ gets larger, the solution gets a longer range, going from exponential to power-law tails. This transition also appears in the energy density. We have also investigated the linear stability to show that the parameter $a$ also modifies the shape of the stability potential, that tends to become a volcano as $a$ gets larger; this impacts the eigenvalues and bound states associated to the Schrödinger-like equation. Previously, in Ref. [14, 36], a similar feature was investigated in a double sine-Gordon model that contains a parameter that plays the same role of $a$. Recently, another mechanism to go from short- to long-range kinks was proposed in a two-field model [46], in which the modification in the tail is induced by one of the fields. In our model, described by the potential (9), the transition from exponential to power-law tails occurs exclusively due to the self-interaction of the field, without extra degrees of freedom.